Find the mean, median, mode, and range for the following list of values:



   13, 18, 13, 14, 13, 16, 14, 21, 13

Grade:Upto college level

1 Answers

Vijay Luxmi Askiitiansexpert
15 years ago

Hi Ritu....

The mean is the usual average, so:

(13 + 18 + 13 + 14 + 13 + 16 + 14 + 21 + 13) ÷ 9 = 15
Note that the mean isn't a value from the original list. This is a common result. You should not assume that your mean will be one of your original numbers.

The median is the middle value, so I'll have to rewrite the list in order:

            13, 13, 13, 13, 14, 14, 16, 18, 2

There are nine numbers in the list, so the middle one will be the (9 + 1) ÷ 2 = 10 ÷ 2 = 5th number:

            13, 13, 13, 13, 14, 14, 16, 18, 21

      So the median is 14.  

The mode is the number that is repeated more often than any other, so 13 is the mode.

The largest value in the list is 21, and the smallest is 13, so the range is 21 – 13 = 8.

            mean: 15
            median: 14
            mode: 13
            range: 8